OpenStudy (anonymous):

helppp ive been stuck on this one for almost a hour What is the value of x? x/6-5=13 @amistre64 @ganeshie8 @johnweldon1993 @undeadknight26

OpenStudy (igreen):

Is it \((\dfrac{x}{6} - 5 = 13)\) or \((\dfrac{x}{6-5} = 13)\)? @igivemedalsandfollow!

OpenStudy (anonymous):

it doesnt let me put tht @iGreen

OpenStudy (igreen):

I'm asking which equation is it? the first one I put or the second one I put?

OpenStudy (anonymous):

oh second?

OpenStudy (igreen):

\(\dfrac{x}{6-5} = 13\) Subtract: \(\dfrac{x}{1} = 13\) \(\dfrac{x}{1} = x\) \(x = 13\)

OpenStudy (anonymous):

so was i right?

OpenStudy (igreen):

Does the equation on your question look like this? \((\dfrac{x}{6} - 5 = 13)\)

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Or does the equation on your question look like this? \((\dfrac{x}{6-5} = 13)\)

OpenStudy (igreen):

Which one? The first or the second?

OpenStudy (anonymous):

first

OpenStudy (anonymous):

god see im that dumb

OpenStudy (igreen):

Okay.

OpenStudy (igreen):

\(\dfrac{x}{6} - 5 = 13\) Add 5 to both sides: \(\dfrac{x}{6} = 18\) Now you have to multiply 6 to both sides of the equation. Do you know what the answer is now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Well if you have to multiply 6 to both sides, and 18 is on the other side, what's 6 x 18?

OpenStudy (anonymous):

108

OpenStudy (igreen):

Yes, so that's your answer.

OpenStudy (anonymous):

thanks could u help me with a couple more?